OpenStudy (anonymous):

Find the linearization l(x) of the function at a. F(x)=sinx a=pi/6

OpenStudy (amistre64):

if we have a slope and a point, could we make a line equation with them?

OpenStudy (anonymous):

It doesn't give me a point or slope just that

OpenStudy (amistre64):

oh it s gives it to you ...

OpenStudy (amistre64):

the point is a,f(a) and the slope is defined as f'(a)

OpenStudy (amistre64):

y - f(a) = m(x-a) y - f(a) = f'(a) (x-a)

OpenStudy (anonymous):

So would I take the derivative of sin then plug in pi/6 which is1/2

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

So my answer would just be f'(x)=cos x=cos1/2

OpenStudy (amistre64):

y = mx - ma + f(a) a = pi/6 f(a) = sin(pi/6) m=f'(a) = cos(pi/6) y = cos(pi/6)x -pi/6 cos(pi/6) + sin(pi/6) simplify as wanted

OpenStudy (anonymous):

Oh ok I see it now I'm overthinking this thank you

OpenStudy (amistre64):

good luck :)